OpenStudy (anonymous):

m,n,p,q are integers; r=m/n=p/q; q,n>0. prove using only field axioms (A-axioms, M-axioms, and distributive law): (b^m)^(1/n)=(b^p)^(1/q).

OpenStudy (anonymous):

\[m, n, p, q \in \mathbb{Z}; n, q>0; r=m/n=p/q\]\[Prove:(b^n)^{1/m}=(b^p)^{1/q}\]

OpenStudy (chihiroasleaf):

what do you mean by A-axioms and M-axioms? :D